### v2

# DeferredGS: Decoupled and Editable Gaussian Splatting with Deferred Shading

Abstract
Reconstructing and editing 3D objects and scenes both play crucial roles in computer graphics and computer vision. Neural radiance fields (NeRFs) can achieve realistic reconstruction and editing results but suffer from inefficiency in rendering. Gaussian splatting significantly accelerates rendering by rasterizing Gaussian ellipsoids. However, Gaussian splatting utilizes a single Spherical Harmonic (SH) function to model both texture and lighting, limiting independent editing capabilities of these components. Recently, attempts have been made to decouple texture and lighting with the Gaussian splatting representation but may fail to produce plausible geometry and decomposition results on reflective scenes. Additionally, the forward shading technique they employ introduces noticeable blending artifacts during relighting, as the geometry attributes of Gaussians are optimized under the original illumination and may not be suitable for novel lighting conditions. To address these issues, we introduce DeferredGS, a method for decoupling and editing the Gaussian splatting representation using deferred shading. To achieve successful decoupling, we model the illumination with a learnable environment map and define additional attributes such as texture parameters and normal direction on Gaussians, where the normal is distilled from a jointly trained signed distance function. More importantly, we apply deferred shading, resulting in more realistic relighting effects compared to previous methods. Both qualitative and quantitative experiments demonstrate the superior performance of DeferredGS in novel view synthesis and editing tasks.
Index Terms:
Gaussian Splatting, Inverse Rendering, Editing.
1 Introduction
Reconstructing and editing 3D scenes from multi-view images is an important task in computer graphics and computer vision. Recently, Neural Radiance Fields (NeRFs) [1] have shown promising results in novel view synthesis, based on which a number of editing methods [2, 3, 4, 5] are proposed. However, NeRF-based methods require dense sampling of camera rays, making rendering less efficient. Gaussian splatting [6] instead approximates a scene with a set of 3D Gaussians and visualizes it with a rasterization-based renderer to achieve real-time performance. However, despite high-quality and efficient rendering, Gaussian splatting still lacks the ability of editing.
In the original Gaussian splatting representation, the appearance is approximated by a view-related spherical harmonic function which entangles both texture and lighting information, making separate texture and lighting editing impossible. Independent editing requires texture and lighting decomposition so that one part can be manipulated without influencing the other. However, different from 3D representations like textured meshes or NeRFs that have an explicit surface or a continuous implicit field, which can provide a good geometry basis like normal direction for the texture and lighting decomposition, Gaussian splatting models a scene’s geometry with a discrete set of Gaussians, and this discrete representation makes the geometry less tractable, posing challenges for later texture and lighting decoupling.
Figure 1: Given multi-view images, DeferredGS optimizes a Gaussian splatting representation with decoupled geometry, texture, and lighting. With this decoupled representation, DeferredGS not only allows geometry and texture editing, but can also render the input scene (the 3rd column, bottom row) or the edited scene (the last column, bottom row) under a novel illumination.
Recently, in the context of Gaussian splatting, a few methods have made attempts to edit it. Two independent works both named GaussianEditor [7, 8] utilize the 2D segmentation model [9] and text-to-image diffusion models [10] to enable scene editing like object removal/insertion or text-guided modification similar to Instruct-NeRF2NeRF [11]. To enable better-controlled editing on texture and lighting, the works [12, 13, 14, 15] try to disentangle texture and lighting to enable individual editing for both diffuse scenes and reflective scenes. Although obtaining plausible decomposition results, they use forward shading that first calculates shading for each Gaussian and then blends shaded Gaussians, which can cause blending artifacts when the scene is relighted. This is because the geometry attributes of Gaussians like opacity are optimized under the input lighting condition while under novel illuminations shaded Gaussians might be blended in a wrong way (see 3rd to 6th rows in Fig. 5 for some examples). In addition, they learn geometry from the Gaussian splatting representation with the normal direction derived from the rendered depth map so their geometry may contain noise and degenerate on scenes with reflective surfaces (see the “Normal” rows in Fig. 4).
To resolve the aforementioned issues, we propose DeferredGS that decouples a scene’s texture and its surrounding illumination of the Gaussian splatting representation with the deferred shading technique. Given multi-view captures of an input scene, we define a set of Gaussians with extra learnable texture attributes and normal directions attached to model texture and geometry respectively. To enable faithful geometry reconstruction, we distill the normal field from a signed distance function (SDF) onto the Gaussians. We also optimize a learnable environment map to approximate the surrounding lighting conditions. In the rendering process, geometry and texture buffers are first rasterized and shading is calculated at the pixel level with the deferred shading technique to enable faithful decoupling and editing. Despite involving more complex shading computation, DeferredGS still allows real-time rendering ($\sim$30FPS at $800\times 800$ resolution on a single 3090 GPU).
Our contributions are as follows:
  • We introduce DeferredGS, a decoupled Gaussian splatting representation with editable geometry, texture and lighting, where its geometry is enhanced by a normal distillation module.
  • To the best of our knowledge, DeferredGS is the first to apply the deferred shading technique to Gaussian splatting, which alleviates blending artifacts of previous methods.
  • Experiments show that our DeferredGS produces more faithful decomposition and editing results compared to previous methods.
2 Related Work
2.1 Neural Rendering
With the development of recent implicit representations [16, 17, 18] and volume rendering [19], neural rendering has become popular in 3D scene reconstruction. The pioneering work Neural Radiance Field (NeRF) [1] models a sample point’s density and view-dependent color with two neural networks and renders the scene by integrating sample points’ colors along the rays. NeRF can synthesize realistic novel view results but dense sampling and network query make both training and rendering slow. To improve the efficiency, a few hybrid rendering methods bake view-independent features onto regular or irregular geometry proxies. NSVF [20] models 3D scenes with an octree structure and prunes voxels with low-density values so that sample points outside the octree can be ignored in the rendering process. DVGO [21] stores density values and appearance features on two voxel grids and directly optimizes these two voxel grids to reconstruct 3D scenes. TensoRF [22] models a scene with a set of matrices and vector pairs that depict the scene’s geometry and appearance on corresponding planes, which not only accelerates the training speed but also reduces the storage size. Instant-NGP [23] instead utilizes multi-resolution hash encoding as its scene representation and enables training convergence in as little as 5 seconds. MobileNeRF [24] and BakedSDF [25] share a similar idea to first reconstruct the mesh of a scene and attach learnable appearance information onto mesh vertices, which enables real-time rendering on consumer devices. Moving a step forward, Gaussian splatting [6] abandons the neural network and models a 3D scene as a set of 3D Gaussians on which attributes like position, rotation, scaling, opacity, and spherical harmonic coefficients can be tuned. Gaussian splatting uses a rasterization-based renderer which can produce rendered results from a viewpoint in real time. To better reconstruct reflective scenes, GaussianShader [14] proposes to model view-dependent effects separately similar to RefNeRF [26] and Ye et al. [27] further introduce deferred rendering for reflection modeling. Despite the high-quality novel view results, these methods still lack the capability of editing, making appearance change or relighting impossible.
2.2 Neural Surface Reconstruction
Although NeRF and Gaussian splatting can synthesize realistic novel view results, they do not have an explicit surface representation, limiting their application in the traditional graphics pipeline. Therefore, a few methods propose to build the connection between the distance field and the density field in volume rendering. UniSurf [28] models the geometry as an occupancy function and replaces alpha values in volume rendering with occupancy values. NeuS [29] instead treats the geometry as a signed distance field (SDF) and proposes an unbiased and occlusion-aware transformation from SDF to density field. VOLSDF [30] also uses SDF as its geometry representation but is based on a Laplace distribution’s Cumulative Distribution Function. To reconstruct open surfaces, several concurrent works [31, 32, 33] introduce the unsigned distance field (UDF) into the training of NeRF. However, for shiny scenes, these methods may produce unsatisfactory results like concave surfaces. To address this, the works [34, 5, 35] decompose the appearance of a scene into view-independent appearance and view-dependent appearance based on RefNeRF [26] and better reconstruct shiny scenes. In the field of Gaussian splatting, NeuSG [36] introduces a NeuS [29] branch to align Gaussians’ normals with corresponding normal directions in NeuS but requires a long training time (16 hours) compared to the original Gaussian splatting. SuGaR [37] instead introduces a depth self-regularization loss term to constrain Gaussians to lie on the surface.
2.3 Inverse Rendering
The goal of inverse rendering is to recover geometry, material and lighting information from a set of images of a single scene under the same or different illuminations. NeRV [38] is the pioneering work that decomposes geometry and BRDF (Bidirectional Reflectance Distribution) material under a given lighting condition. To reduce dependence on light input, later methods work under unknown light conditions. NeRD [39] and PhySG [40] use Spherical Gaussian (SG) as the light representation and utilize a density field and a signed distance field as their geometry representations. Based on PhySG, InvRender [41] further takes visibility into consideration to better recover material information in occluded regions. NeROIC [42] approximates the illumination with Spherical Harmonic functions and enables training on Internet images under different lighting conditions. NeRFactor [43] represents the lighting with a low-resolution image. To better estimate material from images, it also trains a BRDF autoencoder. NDR [44] and NDRMC [45] use a hybrid geometry representation DMTet [46] to enable efficient inverse rendering. NMF [47] and TensoIR [48] introduce the importance sampling strategy for effective shading calculation. To model reflective objects, NeRO [49] and DE-NeRF [5] express high-frequency lighting as MLP networks. Apart from works concentrating on object-level scenes, there are methods [50, 51, 52] targeting large-scale outdoor scenes. There are also concurrent inverse rendering works based on the Gaussian splatting representation like [12, 13, 15, 14]. Different from them, our method introduces a normal distillation module to enable more accurate normal estimation and alleviate blending artifacts with the deferred shading technique when relighting Gaussians.
3 Method
We propose DeferredGS, a decoupled Gaussian splatting representation that decomposes geometry, texture, and lighting from multi-view images, which enables downstream applications like scene relighting and texture editing. The overview of our method is shown in Fig. 2. First, we briefly introduce the rendering formulation in Gaussian splatting and define the texture parameters we use in rendering (Sec. 3.1). Since directly optimizing everything from scratch can lead to suboptimal results, especially on the geometry side, we introduce a normal field distillation module (Sec. 3.2) that utilizes an SDF network to guide Gaussians to a better surface representation. With geometry set up properly, we bind texture parameters onto the Gaussian splatting representation and perform deferred shading to obtain plausible decomposition (Sec. 3.3). Finally, we demonstrate how this decomposition enables scene editing applications like relighting and texture editing (Sec. 3.4).
Figure 2: Overview of DeferredGS. Given multi-view images, DeferredGS builds a decoupled Gaussian splatting representation where auxiliary texture attributes are defined on each Gaussian. To enable successful geometry and appearance separation, we attach an extra normal direction to each Gaussian and optimize it by distilling the normal field from a jointly trained signed distance function. For texture and lighting decomposition, DeferredGS rasterizes geometry and texture attributes into buffer maps and computes shading at the pixel level under the illumination of a learnable environment map with deferred shading.
3.1 Gaussian Splatting and Texture Definition
Gaussian splatting models a 3D scene by a set of Gaussians with which attributes like position $P$, covariance $\Sigma$, scale $S$, opacity $\alpha$ and Spherical Harmonic coefficients (SH) representing view-dependent appearance are associated. It uses a point-based $\alpha$-blending rendering formulation similar to NeRF [1].
$C_{*}=\sum_{i=1}^{N}*_{i}\alpha_{i}T_{i}$ (1)
where $N$ is the number of sample points on a ray. $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is accumulated transmittance and * can be any attribute of the sample point to be blended such as depth, color and normal direction. Gaussian splatting avoids densely sampling on the ray and efficiently projects 3D Gaussians onto a 2D image plane and approximates 3D scenes by blending view-dependent color represented by SH. Instead of modeling the appearance as a whole with SH, our DeferredGS replaces SH with optimizable texture parameters including diffuse albedo $k_{d}$, roughness $r$, and specular albedo $k_{s}$ as shown in Fig. 2 to enable texture and lighting decoupling.
3.2 Normal Field Distillation
[…]
$L_{nerf}=\sum_{v{\in V}}\left\|C_{c}^{\prime}(v)-C^{t}(v)\right\|+\lambda\sum_{v{\in V}}\sum_{i{=1}}^{M}\left\|\|\nabla_{x_{v,i}}\|-1\right\|_{2}^{2},$ (2)
where $V$ is the number of rays in a batch. $M$ is the number of sample points on a single ray and $C^{t}(v)$ is the corresponding ground truth color. The second term is the Eikonal loss, where $\|\nabla f_{s}({x_{v,i})}\|$ is the spatial norm of the SDF network $f_{s}(x)$’s gradient at the $i$th sample point $x_{v,i}$ on the ray $v$. In order to distill the normal information from Instant-RefNeuS to Gaussian splatting, we define an extra normal direction $n$ on each Gaussian as shown in Fig. 2 and render the normal $C_{n}$ of Gaussian splatting for a pixel using Eqn. (1). Then we distill the normal field Gaussian splatting by minimizing the difference between $C_{n}$ and the corresponding rendered normal in Instant-RefNeuS:
$L_{nd}=\sum_{v\in V}\left\|1-C_{n}(v)\cdot C_{n}^{\prime}(v)\right\|=\sum_{v\in V}\left\|1-C_{n}(v)\cdot C_{\nabla x_{v,i}}\right\|$ (3)
where $\cdot$ is the dot product operator.
3.3 Deferred Shading in Gaussian Splatting
For shading calculation, we follow the rendering equation [53]:
[…]
where $L_{o}(\omega_{o})$ and $L_{i}(\omega_{i})$ represent outgoing lighting in direction $\omega_{o}$ and incident lighting from direction $\omega_{i}$ respectively. $f(\omega_{i},\omega_{o})$ is the BRDF at a point which can be parameterized by the Disney shading model [54]:
$f(\omega_{i},\omega_{o})=\frac{k_{d}}{\pi}+\frac{DFG}{4(\omega_{i}\cdot n)(\omega_{o}\cdot n)}$ (5)
where $D,F,G$ correspond to normal distribution function, Fresnel term, and geometry term. Their detailed calculations can be found in [55].
We model the scene’s environment lighting with a $6\times 512\times 512$ High Dynamic Range (HDR) cube map and utilize the SplitSum [56] approximation which separates the integrals of lighting and BRDF to enable efficient shading calculation. T he diffuse color $c_{diff}$:
[…]
which is only dependent on the normal direction $n$ and can be pre-computed and stored in a 2D texture.
For the specular color $c_{spec}$:
$c_{spec}\approx Int_{light}\cdot Int_{BRDF}$ (7)
[…]
where $Int_{light}$ represents the integral of incident light with the normal distribution function $D$. Given a fixed environment map, $Int_{light}$ is only related to the roughness value $r$ and therefore can be pre-computed and stored in a mipmap where each mip level corresponds to a fixed roughness value. $Int_{BRDF}$ is the integral of specular BRDF under a uniform white environment lighting. It is determined by the roughness value in the BRDF and the dot product between incident light direction and normal direction $\omega_{i}\cdot n$. Again it can be pre-computed and stored in a 2D texture and efficiently looked up in rendering. The final shaded color is $c_{g}=c_{diff}+c_{spec}$.
A straightforward way to obtain a decoupled Gaussian splatting representation is performing forward shading for each Gaussian to get its color $c_{i}$ and rasterizing Gaussians with Eqn. (1) similar to previous works [41, 43, 15, 14]. Though obtaining plausible decoupled results, we observe notable artifacts when rendering the decoupled Gaussians under a novel lighting condition (see “F.S.” columns in Fig. 8 and “RGS” and “GShader” columns in Fig. 5). This is because even though Eqn. (3) constrains the rasterized normal to be close to that of Instant-RefNeuS at pixel level, the rasterized normal is still blended from multiple Gaussians’ normals. The geometry attributes like position, covariance and opacity may overfit the given lighting condition to produce proper blended normal and shaded color and it does not guarantee that each Gaussian’s individual normal or shaded color is correct. Thus when changing the lighting condition, the trained geometry attributes may not work for the shaded colors under a new lighting condition and blend them in a wrong way.
To resolve this issue, we introduce the deferred shading technique into the rendering of Gaussians. That is to say, we first rasterize components like position $P$, normal $n$, diffuse albedo $k_{d}$, roughness $r$, specular albedo $k_{s}$, and opacity $\alpha$ into 2D pixels $C_{P},C_{n},C_{k_{d}},C_{r},C_{k_{s}}$, and $C_{\alpha}$ with Eqn. (1). By aggregating all pixels, we get corresponding 2D maps $I_{P},I_{n},I_{k_{d}}$, $I_{r},I_{k_{s}}$, and $I_{\alpha}$. Then we calculate the shaded color $C$ for a pixel with Eqn. (4). Note that since computing shading for every single pixel in the image is time-consuming, we only perform shading on pixels with $C_{\alpha}$ bigger than 0.5. Overall, we optimize the following loss:
$L=$ $L_{nerf}+\lambda_{nd}L_{nd}+\lambda_{L1}L_{L1}+\lambda_{ssim}L_{ssim}+$ (8)
$\lambda_{mask}L_{mask}+\lambda_{TV}L_{TV}$
where $L_{L1}$ and $L_{ssim}$ are the L1 loss and D-SSIM loss between the rasterized image and the ground truth image same as those in [6]. $L_{mask}=\|I_{\alpha}-I_{m}\|_{1}$ is the mask loss that encourages Gaussians to lie in the foreground and $I_{m}$ is the ground truth mask image. $L_{TV}$ is the total variation loss applied on images $I_{k_{d}},I_{k_{s}},I_{k_{s}}$ to encourage smooth texture estimation.
3.4 Editing Decoupled Gaussians
After optimizing the attributes on Gaussians, we obtain a decoupled Gaussian splatting representation and can render the original scene with edited texture or geometry under a novel illumination.
3.5 Relighting
As we extract the environment map of the original scene, we can replace it by a novel environment map with little effort and render the scene with novel illumination using deferred shading.
3.6 Geometry Editing
For geometry editing, instead of directly operating on Gaussians, we first extract a mesh with Instant-RefNeuS and use it as a geometry proxy to guide the deformation of Gaussians. Specifically, we first deform the mesh using the As-Rigid-As-Possible (ARAP) deformation algorithm [57] which solves for rotations and translations of mesh vertices. Then we find the closest neighbor for each Gaussian on the original mesh surface, whose rotation and translation can be obtained via barycentric mapping (based on the mesh triangle that contains the neighboring point). Finally, we deform each Gaussian by applying the rotation and the translation of its corresponding closest neighbor to Gaussian’s position, covariance matrix, and normal direction. By rasterizing deformed Gaussians, we can render the deformed scene.
3.7 Texture Editing
For texture editing, we follow the editing paradigm in traditional 3D modeling software [58] to paint the scene from a given viewpoint. Our texture editing supports changes on all texture components. We first determine which Gaussians need to be adjusted by considering both whether they lie in the editing mask and the difference between the Gaussian depth and the corresponding depth in the depth map rendered with Eqn. (1). Then we can optimize these Gaussians’ texture attributes * by:
${\arg}\min_{*}\|I_{*}^{i}-I_{e}^{i}\|,*\in\{k_{d},r,k_{s}\}$ (9)
where $I_{*}^{i}$ denotes a rendered texture map, e.g. diffuse albedo map, from the input editing viewpoint $i$. $I_{e}^{i}$ is the edited texture map from viewpoint $i$ by putting edited content onto $I_{*}^{i}$.
However, this naive solution only considers the input editing viewpoint and may produce blending artifacts when we view the optimized Gaussians from other viewpoints. Therefore, we propose to augment the optimization with random view inputs. That is to say, we generate edited rendered results from multiple random viewpoints by first unprojecting the edited content onto the mesh extracted by Instant-RefNeuS to form a textured mesh and rendering the textured mesh from random viewpoints. So the final editing loss is:
${\arg}\min_{*}\sum_{j\in L}\|I_{*}^{j}-I_{e}^{j}\|,*\in\{k_{d},r,k_{s}\}$ (10)
where set $L$ includes both the input viewpoint and randomly sampled viewpoints.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRFactor TensoIR NDR NeRO GS RGS GShader Ours GT
Figure 3: Novel view synthesis comparisons with NeRFactor [43], TensoIR [48], NDR [44], NeRO [49], Gaussian Splatting (GS) [6], RelightableGaussian (RGS) [12], and GaussianShader (GShader) [14]. The bottom two rows are from the real Stanford ORB dataset [59].
Table I: Quantitative comparison of novel view synthesis results using SSIM, PSNR and LPIPS metrics on NeRF Synthetic, Shiny Blender and Stanford ORB datasets.
Methods
NeRF Synthetic
Shiny Blender
Stanford ORB
PSNR
SSIM
LPIPS
PSNR
SSIM
LPIPS
PSNR
SSIM
LPIPS
NeRFactor 27.86 0.924 0.044 27.04 0.913 0.123 30.01 0.934 0.079
TensoIR 29.52 0.944 0.050 28.01 0.905 0.131 34.16 0.980 0.020
NDR 29.05 0.939 0.081 28.11 0.935 0.076 32.84 0.985 0.014
NeRO 29.03 0.942 0.051 30.96 0.953 0.045 29.25 0.970 0.019
GS 33.30 0.969 0.030 30.35 0.944 0.088 35.77 0.991 0.007
RGS 29.25 0.941 0.042 28.69 0.930 0.081 36.65 0.992 0.006
GShader 30.84 0.933 0.067 30.72 0.951 0.069 35.43 0.990 0.008
Ours 32.29 0.971 0.039 32.61 0.968 0.041 37.93 0.995 0.006
4 Results and Evaluations
4.1 Datasets and Metrics
We conduct qualitative and quantitative experiments on two synthetic datasets, NeRF Synthetic [1] and Shiny Blender [26] and one real dataset, the Stanford ORB dataset  [59]. For all datasets, we evaluate DeferredGS and baseline methods on novel view synthesis, decomposition and relighting tasks. For quantitative comparisons, we use PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity) [60] and LPIPS (Learned Perceptual Image Patch Similarity) [61] metrics to compare the similarity between images. To evaluate the quality of the estimated normals, we compare the estimated normal images with the ground truth with Mean Squared Error (MSE).
4.2 Implementation Details
We train the decoupled Gaussian splatting representation on a single 3090 GPU with 24GB memory. The training process takes 3-4 hours depending on the scenes’ complexity and the input images’ resolution. DeferredGS allows $\sim$30FPS novel view synthesis at $800\times 800$ resolution on a single 3090 GPU. For hyperparameter settings, in Eqn. (2), $\lambda_{1}$ is set to $0.1$. In Eqn. (8), $\lambda_{nd},\lambda_{L1},\lambda_{ssim},\lambda_{mask},\lambda_{TV}$ are set to $0.5,0.8,0.2,0.5,0.001$ in all our examples.
4.3 Novel View Synthesis
We present novel view synthesis results on different cases’ test splits in Fig. 3 and compare our method with NeRFactor [43], TensoIR [48], NDR [44], NeRO [49], Gaussian Splatting [6], RelightableGaussian [12] and GuassianShader [14]. NeRFactor models the geometry with a density function and utilizes a low-resolution environment map as its lighting representation, which leads to less convincing geometry (first row in Fig. 3) and blurry reconstruction results. TensoIR is unable to synthesize appearance details with its smooth lighting representation approximated by Spherical Gaussian functions. NDR uses deformable tetrahedra to approximate the geometry which may cause irregular geometry and noisy novel view synthesis results. NeRO produces blurry appearance (second and fourth rows) and may reconstruct wrong geometry (first row). Gaussian Splatting models the texture and lighting with a single spherical harmonic function which may fail to generalize to novel views and produce wrong reflections (last row in Fig. 3). RelightableGaussian and GaussianShader separately model texture and lighting but learn geometry from scratch, which causes blurry reconstruction. As for quantitative comparisons, we compare novel view reconstruction results by different methods with the ground truth image using PSNR, SSIM and LPIPS metrics in Table I. Our method achieves comparable results with baseline methods on the NeRF synthetic dataset and outperforms other baselines on the shiny blender dataset.
Normal
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Albedo
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Normal
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Albedo
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Normal
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Albedo
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Normal
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Albedo
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Normal
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Albedo
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
NeRFactor TensoIR NDR NeRO RGS GShader Ours GT
Figure 4: Decomposed result comparisons with NeRFactor [43], TensoIR [48], NDR [44], NeRO [49], RelightableGaussian (RGS) [12], and GaussianShader (GShader) [14]. In every two rows, we show decomposed normal and diffuse albedo components by different methods and compare them with the ground truth. Note that GaussianShader only decomposes diffuse color instead of diffuse albedo so its diffuse albedo results are unavailable.
Table II: Quantitative comparison of decomposed diffuse albedos and normals by different methods. Diffuse albedo is evaluated with PSNR, SSIM and LPIPS metrics while estimated normals are evaluated with MSE. Note that GaussianShader (GShader) [14] only decomposes diffuse color instead of diffuse albedo so its albedo results are unavailable.
Methods
NeRF Synthetic
Shiny Blender
Stanford ORB
PSNR
SSIM
LPIPS
MSE
PSNR
SSIM
LPIPS
MSE
PSNR
SSIM
LPIPS
MSE
NeRFactor 19.01 0.871 0.086 0.098 18.28 0.854 0.106 0.132 28.62 0.958 0.046 0.067
TensoIR 20.29 0.871 0.103 0.056 20.43 0.879 0.194 0.068 29.11 0.969 0.026 0.029
NDR 18.34 0.859 0.116 0.064 19.66 0.865 0.192 0.058 30.67 0.971 0.023 0.043
NeRO 19.69 0.866 0.137 0.135 18.31 0.867 0.187 0.037 24.70 0.948 0.045 0.144
RGS 18.13 0.862 0.115 0.072 18.10 0.847 0.198 0.065 29.33 0.979 0.025 0.049
GShader - - - 0.115 - - - 0.093 - - - 0.037
w/o NFD 16.25 0.821 0.138 0.143 16.58 0.839 0.189 0.102 29.04 0.964 0.039 0.081
Ours 19.14 0.878 0.109 0.053 20.65 0.883 0.188 0.037 31.99 0.977 0.021 0.024
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
Input NeRFactor TensoIR NDR NeRO RGS GShader Ours GT
Figure 5: Relighting comparisons with NeRFactor [43], TensoIR [48], NDR [44], NeRO [49], RelightableGaussian (RGS) [12], and GaussianShader (GShader) [14]. In the first column, we show the input scene and target environment map. Relighting results by different methods and the ground truth are in other columns.
Table III: Quantitative comparison of relighting results using SSIM, PSNR, and LPIPS metrics. Results are averaged over ten different viewpoints with eight different environment maps on NeRF Synthetic and Shiny Blender datasets. For the Stanford ORB dataset, relighting results are evaluated on the provided 20 image-envmap pairs.
Methods
NeRF Synthetic
Shiny Blender
Stanford ORB
PSNR
SSIM
LPIPS
PSNR
SSIM
LPIPS
PSNR
SSIM
LPIPS
NeRFactor 18.40 0.868 0.094 19.79 0.901 0.086 30.03 0.970 0.028
TensoIR 20.25 0.916 0.071 19.17 0.887 0.150 29.60 0.969 0.026
NDR 19.89 0.899 0.066 20.70 0.921 0.063 30.41 0.968 0.023
NeRO 20.09 0.915 0.067 21.86 0.929 0.059 29.67 0.966 0.028
RGS 18.27 0.859 0.073 18.03 0.832 0.155 29.86 0.975 0.018
GShader 18.12 0.838 0.145 20.35 0.917 0.042 30.72 0.976 0.019
Forward Shading 18.21 0.833 0.070 20.05 0.914 0.053 29.94 0.970 0.024
Ours 19.24 0.904 0.057 22.24 0.939 0.049 31.11 0.978 0.017
4.4 Decomposition
To evaluate the decoupling ability of our method, we render optimized diffuse albedo and normal images and compare them with the corresponding ground truth images. Baseline methods include NeRFactor [43], TensoIR [48], NDR [44], NeRO [49], RelightableGaussian [12], and GuassianShader [14] that model texture and lighting separately. As shown in Fig. 4, the smooth or low-resolution lighting representation in NeRFactor and TensoIR may cause incorrect geometry or texture estimation. NDR still suffers from its tetrahedral representation and produces an irregular normal estimation and wrong texture estimation. NeRO produces faithful geometry but less correct texture estimation (first to fourth rows). For RelightableGaussian and GaussianShader, the normals are directly learned from the Gaussian splatting representation and can be rough on shiny scenes, leading to wrong textures. Note that GaussianShader decomposes diffuse color instead of diffuse albedo. Overall our method produces smooth normal estimation without losing geometric details and the extracted textures have less lighting baked in. We also report the PSNR, SSIM and LPIPS metrics for the extracted diffuse albedo components and the MSE value for the normal image in Table II and our method comes on top. Note that before evaluating the albedo component, we compute the scales between estimated normals and the ground truth normals following NeRFactor [43] except for glossy scenes (‘materials’ and ‘mic’ in the NeRF Synthetic dataset and ‘ball’ and ‘toaster’ in the Shiny Blender dataset) since the scales would be all 0 and result in infinite PSNRs for all methods.
4.5 Editing
We show relighting results and compare with NeRFactor [43], TensoIR [48], NDR [44], NeRO [49], RelightableGaussian [12], and GuassianShader  [14] in Fig. 5. Limited by the lighting representations, NeRFactor and TensoIR are unable to render high-frequency reflections. The irregular geometry of NDR makes the scene shaded in a wrong way. NeRO can produce realistic reflection but its material estimation is less accurate, leading to less faithful relighting (third to fourth rows). In addition, it directly uses the mesh extracted by an SDF network for relighting, which may produce level-set stripe-like artifacts (better viewed in video) or miss geometry details like the metal frame on the side window in the car example. GaussianShader and RelightableGaussian produce less faithful normal and texture estimation in the decomposition process and the forward rendering technique they apply causes blending artifacts even if the normal estimation is plausible (zoom-in views of the car example in Fig. 5). In contrast, the relighting results produced by our method are more realistic. This is because the geometry of DeferredGS distilled from a signed distance function, which works as a good basis for the geometry, texture and lighting decomposition. In addition, DeferredGS uses the deferred shading, which avoids blending artifacts when rendering Gaussians in a new illumination. Quantitative results on the relighting task are reported in Table III and our method achieves comparable results with NeRO [49]. We show geometry and texture editing results in Fig. 6. For texture editing results in the last two rows, the editing is performed on the diffuse component.
Geometry
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Geometry
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Texture
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Texture
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Texture
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input Edit View #1 View #2 View #3
Figure 6: Geometry editing and texture editing results. For each row, we show input and edited geometry/image in the first two columns, and in the remaining three columns we show edited Gaussians rendered from three different viewpoints.
4.6 Ablation
In this subsection, we conduct several ablation studies to prove the effectiveness of our design choices.
4.7 Normal Field Distillation
To produce a plausible geometry for the decomposition problem, we distill the normal direction from a signed distance function by encouraging the rendered normal of Gaussians to be consistent with the normal of a signed distance function. Here, we generate the decomposed results without normal field distillation (w/o NFD) and compare them with the full pipeline qualitatively in Fig. 7. We also report numerical results of decomposed results in Table II and the decomposition quality sees a drop without normal field distillation.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Input
(b) w/o NFD
(c) w/ NFD
(d) w/o NFD
(e) w/ NFD
Figure 7: Qualitative comparisons between decomposed normals and diffuse albedos with normal field distillation (w/ NFD) and without normal field distillation (w/o NFD).
4.8 Deferred Shading
In the decomposition and relighting process, we utilize the deferred shading technique instead of forward shading commonly used in previous methods. Here we ablate its influence on the relighting results in Fig. 8. Even though the geometry is enhanced by the normal field distillation, we can still observe notable Gaussian shape—like blending artifacts when we apply the forward shading technique. By contrast, the deferred shading technique we use can successfully resolve this issue and produce smooth and realistic relighting results. We also evaluate it quantitatively in Table III and the relighting quality drops when it is not applied.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Input
(b) F.S.
(c) D.S.
(d) Input
(e) F.S.
(f) D.S.
Figure 8: Qualitative comparisons between relighting results with forward shading (F.S.) and with deferred shading (D.S.). The relit scene contains blending artifacts when we use forward shading while deferred shading can avoid them. The bottom two rows are from the real Stanford ORB dataset [59].
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Input
(b) Edit
(c) w/o R.V.
(d) w/ R.V.
(e) w/o R.V.
(f) w/ R.V.
Figure 9: Qualitative comparisons between texture editing results with rendered results from random viewpoints (w/ R.V.) and without (w/o R.V.). The rendered results of edited gaussians splatting from novel viewpoints can exhibit inconsistency without inputting rendered results from random viewpoints.
4.9 Random View Inputs in Texture Editing
As mentioned in Sec. 3.7, we perform texture editing on the Gaussian splatting representation by optimizing Gaussians’ attributes to fit the rendered results to both the input edited image and randomly rendered images from different viewpoints. Here we ablate how randomly generated edited results influence the final rendered results of edited Gaussians in Fig. 9. It shows that optimizing Gaussians’ attributes only from the input edited viewpoint can result in overfitting and exhibit inconsistent novel views.
(a) Input
(b) Diffuse Albedo
(c) Edit
(d) Novel View
Figure 10: Limitations. DeferredGS bakes shadow into texture on scenes containing shadow ((a)-(b)) and may result in noises after texture editing ((c)-(d)).
5 Conclusion
In this paper, we present DeferredGS, a Gaussian splatting representation with editable geometry, texture and lighting. To make this happen, we define optimizable texture attributes on Gaussians and an environment map for the physically based rendering process. To ensure proper geometry reconstruction, DeferredGS distills the normal field from previous multi-view reconstruction methods onto Gaussians’ learnable normal attributes. Notably, we propose to use the deferred rendering technique to enable realistic relighting. However, as shown in Fig. 10 our method may produce wrong decoupled results on scenes with shadow. For future directions, it is promising to resolve the ambiguity by capturing scenes under multiple lighting conditions and introducing visibility into the rendering process like [51, 52]. On the editing side, the optimized Gaussians after texture editing may contain noise since Gaussian splatting is a global representation where a pixel is influenced by multiple Gaussians. Adaptive cloning and splitting in the edited region might produce better editing results.